OpenStudy (anonymous):

Find the exact solution to the equation. 6-log9(x+10) =5 need help :O

ganeshie8 (ganeshie8):

start with isolating log term

OpenStudy (anonymous):

6-xlog9+10log9 =5?

ganeshie8 (ganeshie8):

nope. there is no log property like that.. what i meant was : 6-log9(x+10) =5 subtract 6 both sides => -log9(x+10) = -1 multiply -1 both sides => log9(x+10) = 1

ganeshie8 (ganeshie8):

apply the rule : \(\huge \log_a b = c => b = a^c\)

ganeshie8 (ganeshie8):

log9(x+10) = 1 => x+10 = 9^1 => x+ 10 = 9 x = ?

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

oh wow i get it now thx bro :D